### v1

# Wireless Network Topology Inference: A Markov Chains Approach

Abstract
In this work, we address the problem of inferring the topology of a wireless network using limited observational data. Specifically, we assume that we can detect when a node is transmitting, but no further information regarding the transmission is available. We propose a novel network estimation procedure grounded in the following abstract problem: estimating the parameters of a finite discrete-time Markov chain by observing, at each time step, which states are visited by multiple “anonymous” copies of the chain. We develop a consistent estimator that approximates the transition matrix of the chain in the operator norm, with the number of required samples scaling roughly linearly with the size of the state space. Applying this estimation procedure to wireless networks, our numerical experiments demonstrate that the proposed method accurately infers network topology across a wide range of parameters, consistently outperforming transfer entropy, particularly under conditions of high network congestion.
1 Introduction
Inferring the topology of a network from limited information has numerous applications across various fields, including the discovery of gene interactions in computational biology, the analysis of climate dynamics in environmental science, and the investigation of functional brain structure in neuroscience [net_inf_survey]. Additionally, in defense and security contexts, understanding the structure of an adversary’s network can provide valuable insights for decision-making.
In many practical scenarios, the available information about a network often consists of time series data associated with each node. Specifically, in wireless networks, radio frequency sensors deployed in the environment may allow us to detect when each node is transmitting, but not the recipient of the transmission. The objective is then to infer whether pairs of nodes are connected by exploiting correlations in their corresponding time series.
Previous work, particularly in the context of wireless networks [TR13, LC17, TG20, liu2022topology], has focused on detecting links by computing causality measures between time series, such as Granger causality [granger69] and transfer entropy [schreiber00]. While these approaches have demonstrated reasonable accuracy, particularly in simulations, they remain poorly understood. For example, it is unclear in what ways the inferred network approximates the actual network structure.
In this work, we introduce a novel paradigm for network inference. Rather than focusing on local link detection and subsequently attempting to infer global network information by assembling local estimates, we aim to construct a global approximation that better captures the overall structure of the network.
Our goal is to develop a procedure that provably approximates the observed network under a well-defined, albeit simplified, model of network dynamics.
1.1 Our Contribution
We address the problem of estimating the topology of a wireless network from limited information. Specifically, we monitor a network of wireless devices, assuming we can detect when a device is transmitting, but not the target of the transmission. Such data could be collected via a network of radio frequency sensors.
To develop our algorithm, we adopt a Markovian assumption regarding the network’s dynamics: we assume that the future state of the network depends only on its current state and is independent of its past states. While this assumption is not strictly valid for communication networks, we argue that it simplifies the model sufficiently to allow rigorous analysis while still capturing the essential behaviour of wireless networks.
More precisely, we consider the following abstract model: we observe the trajectories of $k\geq 1$ independent discrete-time Markov chains on a finite state space $V$, which represents the set of nodes in the network. Each chain evolves according to a common transition matrix $P$. We assume that we can detect whether a chain is visiting a node at time $t$, but not which specific chain is present. This models the scenario in wireless networks where only node transmissions are observable at time $t$. In this context, $P(u,v)$ denotes the probability that node $v$ will transmit at time $t+1$, given that node $u$ was transmitting at time $t$.
Our primary theoretical contribution is the development of a consistent estimator. We prove that by observing the trajectories of the chains for a sufficient duration, we can construct an estimated transition matrix $\hat{P}$ that approximates the true matrix $P$ arbitrarily well in the operator norm. Crucially, this approximation in operator norm provides insights into the global structure of the network, such as the presence of clusters and bottlenecks [STsparse].
We apply this estimation procedure to wireless networks simulated using the ns-3 network simulator [ns3]. Our experimental results show that the proposed method is competitive with state-of-the-art techniques. Notably, we demonstrate that the estimated topology closely approximates the true topology for various networks with differing structures and sizes. Even in cases where the number of observed transmissions is too small for most links to be detected, our procedure successfully recovers structural information about the network, such as the presence of bottlenecks.
1.2 Related Work
From a theoretical standpoint, our work relates to the broader problem of estimating Markov chain parameters from observed trajectories, an area that has garnered significant attention in recent years [hsu19, Wolf19, WolfKont24]. Our algorithm builds upon techniques developed by Wolfer and Kontorovich [Wolf19]. However, prior work typically focuses on scenarios where only a single Markov chain is observed, which simplifies the estimation task but is less applicable to modelling topology inference in wireless networks. In the single-chain case, observing consecutive transmissions at nodes $u$ and $v$ directly implies the existence of a link between them, a property that does not hold in our model.
There is a significant body of literature on wireless network topology inference (see, e.g.,  [TR13, LC17, salmond19, TG20, mehrotra2021minimax, liu2022topology, du2023network]). However, to the best of our knowledge, our work is the first to provide theoretical guarantees for the estimation procedure, albeit under a simplified model of network dynamics. While this model may not fully capture the complexities of real wireless networks, we believe it encapsulates the key challenges inherent to the inference task.
Finally, given the novelty of our approach, this work represents an initial step in a new line of research on wireless network inference. Further investigation is required to fully assess the strengths and limitations of the methodology. We discuss potential directions for future research in the conclusion section.
2 Problem setup
Let $G=(V,E)$ be a graph representing the wireless network we would like to infer: $V=\{1,\dots,n\}$ represents the set of nodes of the network, while $E\subseteq V\times V$ represents the set of edges (links). We assume $V$ is known and our goal is to infer $E$.
We discretise the time into $T$ intervals. For any time interval $t\in\{1,\dots,T\}$ and any node $u\in V$, we observe if $u$ has emitted a signal during the time interval $t$. We collect this information in a time series matrix $TS\in\{0,1\}^{n\times T}$ such that, for any $u\in V$ and $t\in\{1,\dots,T\}$,
$TS(u,t)=\begin{cases}1&\text{if $u$ is transmitting during time interval $t$}\\ 0&\text{otherwise}.\end{cases}$
Our main task is to infer the set of edges $E$ given the time series matrix $TS$.
To formally model network dynamics, we consider $k$ independent Markov chains $X^{(1)},X^{(2)},\dots,X^{(k)}$ on a finite state space $V$ (the set of nodes of our network) and with the same transition matrix $P$. We denote with $X^{(i)}_{t}$ the position of the $i$-th chain at time $t$. In particular, $X^{(i)}_{t}=u$ means that, at time $t$, node $u$ is transmitting a message to another node. $P(u,v)$ represents the probability that $v$ transmits a message at time $t+1$, given that a message has been transmitted by $u$ at time $t$. Our aim is to infer the transition matrix $P$: this will allows us to infer information not only about the topology of the network, but also information about the role of certain nodes in the network.
For technical reasons, we assume $P$ has a unique stationary distribution $\pi$, i.e., $\pi P=\pi$. This is a reasonable assumption in our context since it essentially means the network is (strongly) connected. Furthermore, we assume the Markov chains are stationary, i.e., $X^{(i)}_{0}\sim\pi$ for all $1\leq i\leq k$. This technical assumption is made to simplify our analysis, but is not necessary and could be dropped with additional technical work (without modifying the proposed algorithm).
We assume $P$ and $\pi$ are unknown to us. Indeed, our goal is to estimate $P$. What we are able to observe is, for each time step $t=0,\dots,T$, how many chains are currently visiting a given vertex $v\in V$. In other words, for any $t=1,\dots,T$ and $v\in V$, we can observe the quantity $S_{t,v}=|\{i\colon X^{(i)}_{t}=v\}|$ for any $t\in\mathbb{N}_{+},v\in V$. We are not able, however, to observe if a specific chain is visiting $v$ at time $t$.
3 Estimation procedure
We are now ready to state our inference procedure. Let $u,v\in V$. We use $[n]$ to denote $\{1,\dots,n\}$. We define the following quantities.
$N(u,v)$ =|$=\left|\{i,j\in[k],t\in[T]\colon X^{(i)}_{t-1}=u,X^{(j)}_{t}=v\}\right|,$
$N(u)$ =|$=\left|\{i\in[k],0\leq t<T\colon X^{(i)}_{t}=u\}\right|.$
In other words, $N(u,v)$ represents the number of times we have observed $u$’s transmission being immediately followed by $v$’s, while $N(u)$ counts the total number of times $u$ is transmitting. Notice that $N(u)=\sum_{v\in V}N(u,v)$. Moreover, these quantities can be computed having access only to $S_{t,v}$ for $0\leq t\leq T,v\in V$.
Let also $M\in\mathbb{R}^{n\times n}$ be defined as
$M(u,v)=\frac{N(u,v)}{N(u)}.$
Essentially, $M(u,v)$ represents the proportion of $u$’s transmissions that are immediately followed by $v$’s.
Let $\widehat{\Pi}$ be the $n\times n$ matrix whose entries are equal to $\widehat{\Pi}(u,v)=N(v)/(kT)$ for any $u,v\in V$. Our matrix estimator $\hat{P}$ is defined as follows:
$\hat{P}=M-(k-1)\widehat{\Pi}.$
The term $(k-1)\widehat{\Pi}$ is a correction necessary because nodes that are visited by the chains more frequently will have larger entries in $M$.
Notice that computing $\hat{P}$ is computationally efficient and can be done in an online manner: we do not have to store the entire time series matrix $TS$, we can just update $N(u,v)$, $N(v)$, and $\hat{P}$ at each time step.
Returning to the wireless network setting, $\hat{P}(u,v)$ can be interpreted as our estimated probability that $v$ will transmit immediately after $u$ has transmitted. While $\hat{P}(u,v)$ is strongly correlated to the likelihood $u$ and $v$ are connected by a link, it is not a measure of how likely a link between $u$ and $v$ exists. To this end, we suggest using an estimator for the Laplacian associated to $P$. For a transition matrix $P$ with stationary distribution $\pi$, the Laplacian $L$ of $P$ is the matrix with entries $L(u,v)=P(u,v)\sqrt{\pi(v)/\pi(u)}$. We recall that any Markov chain can be interpreted as a random walk on a weighted directed graph, where $L(u,v)$ represents the weight on the edge $(u,v)$ [markovmixing]. Therefore, to infer if a link is present or not, we will use the estimated matrix $\hat{L}$, defined as
$\hat{L}(u,v)=\sqrt{\frac{\hat{\pi}(v)}{\hat{\pi}(u)}}\hat{P}(u,v)$
where $\hat{\pi}$ is the leading eigenvector of $\hat{P}$ and an estimate for the stationary distribution $\pi$.
Finally, when we know that links in the observed network are bidirectional, it is natural to consider a symmetrisation of the Laplacian instead:
$\hat{L}_{\text{sym}}(u,v)=\sqrt{\frac{\hat{\pi}(v)}{\hat{\pi}(u)}}\cdot\frac{\hat{P}(u,v)+\hat{P}(v,u)}{2}.$
We conclude this section by underlining that our estimator only records transmissions happening in two consecutive time intervals. This means it’s important to choose an appropriate interval length, otherwise network delays might severely impact our estimator. In our simulations, however, this constraint has not been detrimental to our results. The situation might be more complicated in real world settings. We plan to study how to bypass these potential issues in future work.
4 Theoretical guarantees
To present the main theoretical result of our work, we introduce some additional technical notation. Recall that $n=|V|$ is the number of nodes in our network. We say an event $\mathcal{E}$ holds with high probability if the probability of $\mathcal{E}$ happening is larger than $1-n^{-1}$. We denote with $\sigma_{2}(P)$ the second largest singular value of $P$ and with $\lambda(P)=1-\sigma_{2}(P)^{2}$ the spectral gap of $P$. Given a matrix $M$, we denote with $\|M\|$ its operator norm. We are now ready to state our main theorem.
Theorem 1.
Let $\epsilon\in(0,1/2)$. Define $\pi_{\star}=\min_{u\in V}\pi(u)$ and $\pi^{\star}=\max_{u\in V}\pi(u)$. Assume that
$T>Ck^{3}\epsilon^{-2}\ln(kn)\cdot\lambda(P)^{-1}\frac{\pi^{\star}}{\pi_{\star}^{2}}$
for some large enough universal constant $C>0$. Then, with high probability,
$\|P-(M-(k-1)\widehat{\Pi})\|\leq\epsilon.$
The theorem essentially tells us that $\hat{P}=M-(k-1)\widehat{\Pi}$ is a good estimator for $P$. In particular, by observing the network for a large enough time, we can approximate $P$ arbitrarily well in the operator norm. This approximation preserves important quantities of the network. For example, it approximately preserves its stationary distribution; this tells us how often, on average, nodes are visited [markovmixing], which corresponds how often a node is transmitting in our model. It also preserves its spectral gap, which tells us how fast a Markov chain with transition matrix $P$ converges to its stationary distribution from a worst case starting point [MonteTetali]. Moreover, it preserves important connectivity properties of the networks, such as its “cut-structure”, that is, the structure of clusters and bottlenecks in the network [PSZ, STsparse].
What this approximation doesn’t necessarily preserve are the individual links in the network. For example, consider a transition matrix of a random walk on a complete graph of $n$ vertices. Now consider an Erdos-Renyi random graph $G(n,p)$: that is, a graph of $n$ vertices where we connect any two vertices with probability $p$. As long as $p\gg log(n)/n$, it is well-known that the transition matrix of a random walk on $G(n,p)$ well approximates (in the operator norm) the transition matrix of the complete graph [STsparse]. However, the $G(n,p)$ graph might have only $O(n\log(n))$ edges compared to the $\Theta(n^{2})$ edges of the complete graph, i.e., only a small fraction of the edges are preserved. Indeed, it is possible to construct examples where Theorem 1 guarantees a good approximation with a number of samples that is much smaller than the number of edges in the graph: certain links wouldn’t have been used at all and, therefore, would not be detectable.
We now discuss the sample complexity of our procedure. The number of samples (or time steps) required depends on several parameters. In particular, it is inversely linear in the spectral gap $\lambda(P)$. The spectral gap is a crucial parameter in the analysis of Markov chains: its reciprocal, $1/\lambda(P)$ provides a bound on the mixing time of $P$, i.e., the time it takes for a Markov chain with transition matrix $P$ to converge to the stationary distribution from a worst case starting point [MonteTetali]. A long as $P$ is irreducible, as in our case, $\lambda(P)>0$. We believe that, in cases where $P$ models a communication network, $1/\lambda(P)$ is relatively small. Furthermore, for natural transition matrices $P$, $1/\pi_{\star}$ is proportional to $n$, the size of the network. A larger value of $1/\pi_{\star}$ would mean, in our model, that there are nodes that transmits very rarely: to approximate $P$, we necessarily need to collect enough information about these nodes, increasing the time required. Finally, $\frac{\pi^{\star}}{\pi_{\star}}$ can be interpreted as the ratio between the largest and the smallest frequencies of transmission of a node: if nodes transmit at roughly the same frequency, this parameter becomes order of 1.
In summary, for many network topologies, as long as the number of concurrent transmissions $k$ is not too large, Theorem 1 asserts that a number of samples proportional to the size of the network suffices to well approximate $P$.
The proof of Theorem 1, deferred to Section 7, is inspired by the techniques of [Wolf19]. In particular, it uses matrix concentration inequalities for Markov chains [matrixMarkovBernstein] and martingales [Tropp].
We end this section by briefly discussing the validity of our Markovian model for capturing our task of estimating the topology of a wireless network. Even though communication networks do not satisfy a Markovian assumption, we hope that, on average, the behaviour of such networks can be approximated well enough by a Markovian model. Besides this consideration, however, for our model to capture the behaviour of wireless networks, an additional assumption must be considered. Since multiple chains can concurrently visit a node at any time, our estimation procedure requires to observe the number of chains visiting a node $u$ at a time $t$. In our modelling, however, a chain visiting a node $u$ at time $t$ corresponds to observing node $u$ transmitting a message during the $t$-th time interval. Therefore, $r$ chains visiting $u$ at time $t$ would correspond to $u$ transmitting $r$ messages during the same time interval. This event, besides being very unlikely, cannot even be detected: we assume we can only observe if a node is transmitting in a certain time interval, not how many messages is transmitting. For this reason, our Markovian model can only be realistic if the chance of two or more chains visiting the same node at the same time is very small. Formally, the probability that at a time $t$ two stationary chains visit the same node can be upper bounded by
$\sum_{u\in V}\sum_{1\leq i<j\leq k}\pi(u)={k\choose 2}\sum_{u\in V}\pi(u)^{2}\leq k^{2}\|\pi\|_{2}^{2}.$
Therefore, we assume that $k^{2}\|\pi\|_{2}^{2}\ll 1$ or, equivalently, that
$k\ll\frac{1}{\|\pi\|}.$ (1)
If the stationary distribution of the chain is approximately uniform, we have that $\|\pi\|\approx 1/\sqrt{n}$. In this case, Eq. 1 can be interpreted as $k\ll\sqrt{n}$.
5 Experimental results
We have simulated various wireless network topologies with the network simulator ns-3 [ns3]. Each wireless network is constructed by generating $n$ points on the plane. Each point corresponds to a wireless device and two devices are able to communicate with one another if they are at most 50m apart. Different point sets will induce a different network topology. Propagation speed is constant (the speed of light) and path loss is dependent only on the transmission range: if the receiver is within range, it will receive the transmission at 16dbm, else it will not receive a transmission. We use the IEEE 802.11b protocol with a data rate of 1Mbps. We use the OLSR routing protocol to discover and disseminate edge state information. The Hello and TC intervals are 2 and 5 seconds, respectively. We do not send any packets within the first 30 seconds of the simulation to allow time for devices to calculate the initial shortest path between other nodes in the network. We use UDP as the transport layer protocol and each node is implemented as a UDP client. This choice has been made because UDP does not use acknowledgment messages (ACKs), possibly making link detection more difficult. However, we have also performed some simulations with TCP obtaining very similar results.
We sample 500 pairs of sender/receiver nodes $(i,j)$ uniformly at random (with replacement) and transmit 3 packets of size 100 bytes from $i$ to $j$ at time $30+t$, where $t$ is uniformly distributed in $[0,T]$ for some transmission window length $T>0$.
The time series matrix is constructed by sampling the interval within the simulation between 30 and $30+T$ seconds into intervals of 1.5ms. We remark we haven’t really optimised this number: time-intervals between 1ms and 1.5ms usually work quite well, both for the method outlined in this report and for other methods based on transfer entropy or Granger causality.
Our experiments assume the network is bidirectional, and therefore, we will use our estimate $\hat{L}_{\text{sym}}$ for the symmetrised Laplacian. We will compare our estimate with one derived from transfer entropy. More precisely, we will consider the matrix $M_{\text{TE}}$ where the $(i,j)$-entry is equal to the transfer entropy from the time series associated with $i$ to the time series associated with $j$. To make the comparison even, we will further symmetrised this matrix by taking the average between this matrix and its transpose.
In Fig. 1 we present experimental results for a network topology corresponding to an undirected cycle of $6$ nodes. We can see that both transfer entropy and our methodology works well in recovering links in the network. In Fig. 1(d) we stud the effect of a varying transmission length on network estimation. In particular, we record the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window. More precisely, we say that a link $\{i,j\}$ has been correctly identified by our methodology (resp. transfer entropy) if it appears in the top $m$ entries of the upper triangular part of the $\hat{L}_{\text{sym}}$ (resp. transfer entropy) matrix, where $m$ in the number of links in the network. We observe that a shorter transmission window makes estimation harder since it implies a larger number of nodes will be concurrently transmitting (i.e., a larger parameter $k$ in 1). We notice how our procedure slightly improves upon transfer entropy when the network is congested.
Further experimental results are reported in Figs. 2, 3, 4 and 4. We highlight our methodology consistently outperforms transfer entropy whenever the network is congested. Transfer entropy and our estimate appear very similar when $T\geq 15s$.
Figure 1: Experimental results for a cyclic network topology of six nodes. Figure (a) displays the adjacency matrix of the network. Figure (b) the estimated matrix $\hat{L}_{\text{sym}}$ for a transmission window length $T=15s$. Figure (c) the transfer entropy matrix ($T=15s$). Figure (d) represents the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window. All results are averaged over $50$ independent simulations.
Figure 2: Experimental results for a network topology of eight nodes.
Figure 3: Experimental results for a network topology where nodes are arranged in a three-by-three grid.
Figure 4: Experimental results for a network topology where nodes are arranged in a four-by-two grid.
Figure 5: Experimental results for a network topology of twenty nodes arranged in a path.
6 Conclusion and future work
In this work, we introduced a novel technique for estimating the topological structure of wireless networks. We demonstrated that our approach provably recovers information about network topology under a simplified Markovian model of network dynamics. Additionally, our experimental results showed that the proposed method is competitive with transfer entropy in realistic network simulations, consistently outperforming it under conditions of network congestion.
We believe that, in addition to being competitive with existing methods, our approach also provides valuable insights into the behaviour of traditional techniques. Nevertheless, further research is needed to enhance both the performance and the theoretical understanding of our method. Below, we outline several potential directions for future work.
First, our method currently requires discretisation in the time domain and does not account for variable network delays. To address these limitations, one possible direction would be to consider continuous-time Markov chains. In this framework, a random clock is assigned to each node, and a particle moves from the current node once the clock expires. Since the delays between consecutive packet transmissions would vary, any estimation algorithm would need to account for these fluctuating delays.
Another potential generalisation of our model is to employ branching processes instead of random walks on graphs. In this scenario, particles could appear, move, die, or duplicate according to specific probability distributions. This approach could help address the limitations of our current model, which does not adequately account for varying traffic levels within the network.
Finally, measures such as transfer entropy and Granger causality are often used as features in more sophisticated estimation procedures. For instance, Testi and Giorgetti [TG20] utilise these measures to train a neural network model. It would be interesting to explore the integration of our techniques into machine learning frameworks for network inference.
7 Proofs
We will show that, for $T$ large enough, $M(u,v)\approx P(u,v)-(k-1)\pi(v)$ and $N(v)\approx T\cdot k\cdot\pi(v)$. Therefore, by computing $M$ and $N$, we are able to approximate $P$. We start by showing that $N(v)$ is approximately a multiple of the stationary distribution at $v$. For technical reason, we also need a bound on the following quantities ($u,v\in V$):
$\widetilde{N}(u,v)=\sum_{t=0}^{T-1}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\},$ (2)
$Q(u,v)=\sum_{t=0}^{T-1}\sum_{i,i^{\prime}}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(i^{\prime})}=v\}.$ (3)
We will use the following recent matrix Bernstein inequality for Markov chains [matrixMarkovBernstein].
Theorem 2.
Let $\{X_{i}\}_{i=1}^{\infty}$ be a Markov chain with finite state space $V$, stationary distribution $\pi$, and positive spectral gap $\lambda>0$. Let $f\colon V\to\mathbb{C}^{d\times d}$ such that: (1) $f(u)$ is Hermitian and there exists $A>0$ such that $\|f(u)\|\leq A$ for all $u\in V$; (2) $\mathop{\mathbb{E}}_{#1}[f(v)]=0$; (3) $\|\mathop{\mathbb{E}}_{#1}[f(v)^{2}]\|\leq\sigma^{2}$. Then, for any $\epsilon\in(0,1)$ and $T\in\mathbb{N}$, it holds that
$\mathop{\mathbb{P}}[\left\|\sum_{t=1}^{T}f(X_{t})\right\|>\epsilon]\leq 2d^{2}\exp\left(-\frac{\epsilon^{2}\lambda/72}{T\sigma^{2}+A\epsilon}\right).$
Lemma 3.
Let $\epsilon\in(0,1/2),\delta\in(0,1)$. Let $T>Ck^{2}(\pi^{\star}/\pi_{\star})\ln(kn/\delta)/(\epsilon^{2}\lambda(P)\pi_{\star})$ for some large enough universal constant $C>0$. Let $D_{N}$ be the $n\times n$ diagonal matrix such that $D_{N}(u,u)=N(u)$. Then, with probability at least $1-\delta$, for all $u,v\in V$, the following holds:
$\left|N(v)-kT\pi(v)\right|\leq\epsilon kT\pi(v),$
$\left\|\widetilde{N}-Tk(k-1)\pi\pi^{T}\right\|\leq\epsilon Tk(k-1)\pi_{\star}$
$\left\|Q-Tk(k-1)\pi\pi^{T}-D_{N}\right\|\leq Tk^{2}\pi^{\star}$
Proof.
We first fix $v\in V$ and compute the expectation of $\mathop{\mathbb{E}}[N(v)]$:
$\mathop{\mathbb{E}}[N(v)]$ $=\sum_{i=1}^{k}\sum_{t=0}^{T-1}\mathop{\mathbb{P}}[X^{(j)}_{t}=v]=kT\pi(v),$
where the last equality holds because we assume the chains start from stationarity.
We now have to bound how far $N(v)$ is from its expectation. We do this by applying the scalar version of Theorem 2. Let $f\colon V\to\mathbb{R}$ be defined as $f(u)=\mathbbm{1}\{u=v\}(1-\pi(v))$. Clearly, $|f(u)|\leq 1$ for all $u\in V$, $\mathop{\mathbb{E}}_{#1}[f(u)]=0$ and $\mathop{\mathbb{E}}_{#1}[f(v)^{2}]\leq\pi(v)$. Notice that $N(v)-\mathop{\mathbb{E}}[N(v)]=\sum_{i=1}^{k}\sum_{t=1}^{T}f(X^{(i)}_{t})$. Let us fix $i\in\{1,\dots,k\}$. Then, by Theorem 2 and the assumption on $T$, we have that with probability at most $\delta/(kn)$, $\sum_{t=1}^{T}f(X^{(i)}_{t})\leq\epsilon T\pi(v)$. The first inequality in the statement of the lemma follows by taking a union bound over all $v\in V$ and $i\in\{1,\dots,k\}$.
For the second inequality, we first compute $\mathop{\mathbb{E}}[\widetilde{N}(u,v)]$:
$\mathop{\mathbb{E}}[\widetilde{N}(u,v)]$ $=\sum_{t=0}^{T-1}\sum_{i\neq j}\mathop{\mathbb{E}}[\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\}]$
$=k(k-1)T\pi(u)\pi(v).$
To control the deviation of $\widetilde{N}(u,v)$ from its expectation, fix $i\neq j$ and consider the product chain $(X^{(i)},X^{(j)})$. It has state space $V\times V$ and stationary distribution $\pi((u,v))=\pi(u)\pi(v)$. Moreover, its spectral gap is equal to the spectral gap $\lambda(P)$ of the single chain $X^{(i)}$. Consider $g_{ij}\colon V\times V\to\mathbb{R}^{n\times n}$ defined as follows.
$g_{ij}((u,v))(x,y)=\begin{cases}1-\pi(u)\pi(v)&\text{ if }\{u,v\}=\{x,y\}\\ -\pi(x)\pi(y)&\text{ otherwise.}\end{cases}$
Notice that $\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}]=\sum_{i\neq j}\sum_{t=0}^{T-1}g_{ij}(X_{t}^{(i)},X_{t}^{(j)})$. Clearly, $\|g_{ij}((u,v))\|\leq 2$ for all $(u,v)\in V\times V$. Furthermore, simple calculations show that
$\left\|\mathop{\mathbb{E}}_{#1}[g_{ij}((u,v))^{2}]\right\|\leq 2\pi^{\star}.$
Therefore, with probability at least $\delta/k^{2}$, Theorem 2 implies that
$\left\|\sum_{t=0}^{T-1}g_{ij}(X_{t}^{(i)},X_{t}^{(j)})\right\|\leq Tk\pi_{\star}.$
A union bound over all $1\leq i\neq j\leq k$ yields the second statement.
Finally, the third inequality follows simply from the first two by noticing that $Q(u,v)=\widetilde{N}(u,v)+D_{N}(u,v)$, ∎
We now prove concentration for $M$. We apply similar techniques as in Kontorovich and Wolfer [Wolf19]. We will need the following concentration inequality for matrix martingales, which is due to Tropp [Tropp].
Theorem 4.
Consider a random process $\{Y_{k}\colon k=0,1,2,\dots\}$ such that each $Y_{k}$ is a $d_{1}\times d_{2}$ matrix. Assume this process satisfies the martingale property:
$\mathop{\mathbb{E}}_{#1}[Y_{k}]=Y_{k-1}\quad\text{ for all }k=1,2,\dots,$
where with $\mathop{\mathbb{E}}_{#1}[Y_{k}]$ we denote the conditional expectation of $Y_{k}$ wrt “the past”.
Let $X_{k}=Y_{k}-Y_{k-1}$ for $k=1,2,\dots$, which implies $\{X_{k}\}_{k}$ is a martingale difference sequence for $\{Y_{k}\}_{k}$, i.e.,
$\mathop{\mathbb{E}}_{#1}[X_{k}]=0\quad\text{ for all }k=1,2,\dots.$
Assume the operator norm of $X_{k}$ is bounded by $R$ almost surely:
$\|X_{k}\|\leq R\quad\text{a.s. for all }k=1,2,\dots$
Define $W^{(1)}_{k}=\sum_{j=1}^{k}\mathop{\mathbb{E}}_{#1}[X_{j}^{*}X_{j}]$ and $W^{(2)}_{k}=\sum_{j=1}^{k}\mathop{\mathbb{E}}_{#1}[X_{j}X_{j}^{*}]$ for $k=1,2,\dots$. Further assume their operator norm is bounded by $\sigma^{2}$:
$\max\left(\|W^{(1)}_{k}\|,\|W^{(2)}_{k}\|\right)\leq\sigma^{2}\quad\text{a.s. for all }k=1,2,\dots$
Then, for all $t>0$, it holds that
$\mathop{\mathbb{P}}[\exists k\geq 0\colon\|Y_{k}\|\geq t]\leq(d_{1}+d_{2})\exp\left(-\frac{t^{2}/2}{\sigma^{2}+Rt/3}\right).$
We define the following random variables. For any $u,v\in V$ and $t=1,2,\dots,T$,
$Y_{t}(u,v)=\sum_{i=1}^{k}\mathbbm{1}\{X_{t-1}^{(i)}=u\}\left(\mathbbm{1}\{X_{t}^{(i)}=v\}-P(u,v)\right)$
$+\sum_{1\leq i\neq j\leq k}\mathbbm{1}\{X_{t-1}^{(i)}=u\}\mathbbm{1}\{X_{t}^{(j)}=v\}$
$-\sum_{1\leq i\neq j\leq k}\mathbbm{1}\{X_{t-1}^{(i)}=u\}\sum_{w\in V}\mathbbm{1}\{X_{t-1}^{(j)}=w\}P(w,v).$ (4)
Notice that
$\mathop{\mathbb{E}}_{#1}[\mathbbm{1}\{X_{t-1}^{(i)}=u\}\mathbbm{1}\{X_{t}^{(i)}=v\}]=\mathbbm{1}\{X_{t-1}^{(i)}=u\}P(u,v)$
while, for $i\neq j$,
$\mathop{\mathbb{E}}_{#1}[\mathbbm{1}\{X_{t-1}^{(i)}=u\}\mathbbm{1}\{X_{t}^{(j)}=v\}]=\mathbbm{1}\{X_{t-1}^{(i)}=u\}\sum_{w\in V}\mathbbm{1}\{X_{t-1}^{(j)}=w\}P(w,v)$
Therefore,
$\mathop{\mathbb{E}}_{#1}[Y_{t}(u,v)]=0,$
which implies that $Y_{t}$ is a martingale difference sequence.
Let $\widetilde{N}(u,v)=\sum_{t=0}^{T-1}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\}$. Then, for any $u,v\in V$,
$\sum_{t=1}^{T}Y_{t}(u,v)$ =|$=\left|\{i,j\in[k],t\in[T]\colon X^{(i)}_{t-1}=u,X^{(j)}_{t}=v\}\right|$
|$-\left|\{i\in[k],0\leq t<T\colon X^{(i)}_{t}=u\}\right|P(u,v)$
$-\sum_{w\in V}\sum_{t=0}^{T-1}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=w\}P(w,v)$
$=$ $N(u,v)-N(u)P(u,v)-\sum_{w\in V}\widetilde{N}(u,w)P(w,v).$
Notice the previous equation can be rewritten in matrix form as
$\sum_{t=1}^{T}Y_{t}=N-(D_{N}+\widetilde{N})P.$
Furthermore,
$\mathop{\mathbb{E}}[\widetilde{N}P(u,v)]$ $=\sum_{w\in V}\widetilde{N}(u,w)P(w,v)$
$=Tk(k-1)\pi(u)\sum_{w\in V}\pi(w)P(w,v)$
$=Tk(k-1)\pi(u)\pi(v),$
where the last equality follows because $\pi$ is the stationary distribution of $P$.
This combined with the fact that $\mathop{\mathbb{E}}[N(u)]=Tk\pi(u)$, suggests $\widetilde{M}P(u,v)\approx(k-1)\pi(v)$, where $\widetilde{M}=D_{N}^{-1}\widetilde{N}$. Indeed, our aim is to prove that $M\approx P+(k-1)\Pi$ where $\Pi(u,v)=\pi(v)$.
By Lemma 3, with high probability, $\|\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}]\|<\epsilon/10$. Therefore,
$\|\widetilde{N}P-k(k-1)T\pi\pi^{T}\|$ $<\|\mathop{\mathbb{E}}[\widetilde{N}]P-k(k-1)T\pi\pi^{T}\|+\|(\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}])P\|$
$\leq\|\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}]\|\|P\|$
$\leq\frac{Tk\pi_{\star}\epsilon}{10}.$
Hence,
$\|D_{N}(\widetilde{M}P-(k-1)\Pi)\|<\frac{Tk\pi_{\star}\epsilon}{8}.$ (5)
We will apply Theorem 4 to show that $\|N-(D_{N}+\widetilde{N})P\|$ is small. This, together with Equation (5) will imply the theorem.
We start by obtaining a bound on $\|Y_{t}\|$.
Lemma 5.
Let $Y_{t}$ ($t=1,2,\dots$) be defined as in (4). Then, for any $t$, it holds that
$\|Y_{t}\|\leq\sqrt{2}k^{2}.$
Proof.
We bound $\|Y_{t}\|$ by applying Holder’s inequality for matrix norms: $\|Y_{t}\|\leq\sqrt{\|Y_{t}\|_{1}\|Y_{t}\|_{\infty}}$, where $\|Y_{t}\|_{1}=\max_{v\in V}\sum_{u\in V}|Y_{t}(u,v)|$ and $\|Y_{t}\|_{\infty}=\max_{u\in V}\sum_{v\in V}|Y_{t}(u,v)|$.
We start with a bound on the $\ell_{1}$ norm.
$\|Y_{t}\|_{1}$ $=\max_{v\in V}\sum_{u\in V}|Y_{t}(u,v)|$
$\leq\max_{v\in V}\left\{|\sum_{i\colon X^{(i)}_{t}=v}(1-P(X^{(i)}_{t-1},v))|+k\cdot\sum_{i=1}^{k}|\sum_{j\colon X^{(j)}_{t}=v}(1-P(X^{(j)}_{t-1},v))|\right\}$
$\leq k+(k-1)k=k^{2}.$
For the $\ell_{\infty}$ norm, first notice that
$\sum_{v\in V}\left|\left(\mathbbm{1}\{X_{t}^{(i)}=v\}-P(u,v)\right)\right|$ $=1-P(u,X_{t}^{(i)})+\sum_{w\neq X_{t}^{(i)}}P(u,w)$
$=2(1-P(u,X_{t}^{(i)})).$
Therefore,
$\|Y_{t}\|_{\infty}$ $=\max_{u\in V}\sum_{v\in V}|Y_{t}(u,v)|$
$\leq\max_{u\in V}\left\{2\sum_{i=1}^{k}(1-P(u,X_{t}^{(i)}))+\sum_{i=1}^{k}\cdot|\sum_{j\neq i}2(1-P(X^{(j)}_{t-1},v))|\right\}$
$\leq 2k+2(k-1)k=2k^{2}.$
Hence,
$\|Y_{t}\|\leq\sqrt{\|Y_{t}\|_{1},\|Y_{t}\|_{\infty}}\leq\sqrt{2}k^{2}.$
The following lemma provides a bound on $\sigma^{2}$ as required by Theorem 4. The proof of the lemma is not particularly enlightening but it is relatively involved, since it requires keeping track of different terms arising from the multiplication of $Y_{t}$ with its transpose.
Lemma 6.
Let $Y_{t}$ ($t=1,2,\dots$) be defined as in (4). Let $W^{(1)}_{t}=\sum_{j=1}^{t}\mathop{\mathbb{E}}_{#1}[Y_{j}^{T}Y_{j}]$ and $W^{(2)}_{k}=\sum_{j=1}^{t}\mathop{\mathbb{E}}_{#1}[Y_{j}Y_{j}^{T}]$. Then, with high probability, it holds that
$\|W^{(1)}_{T}\|\leq 8k^{3}T\pi^{\star}.$
and
$\|W^{(2)}_{T}\|\leq 2k^{3}T\pi^{\star}.$
Proof.
We first obtain a bound on the operator norm of $W^{(1)}_{t}$ Let $u,v\in V$. Then,
$Y_{j}^{T}Y_{j}(u,v)=\sum_{w\in V}Y_{j}(w,u)Y_{j}(w,v)$
$=\sum_{w\in V}\sum_{i=1}^{k}\mathbbm{1}\{X_{j-1}^{(i)}=w\}\left(\mathbbm{1}\{X_{j}^{(i)}=u\}-P(w,u)\right)$
$+\sum_{w\in V}\sum_{1\leq i\neq\ell\leq k}\mathbbm{1}\{X_{j-1}^{(i)}=w\}\left(\mathbbm{1}\{X_{j}^{(\ell)}=u\}-P(X_{j-1}^{(\ell)},u)\right)$
$\quad\cdot\left(\sum_{i^{\prime}=1}^{k}\mathbbm{1}\{X_{j-1}^{(i^{\prime})}=w\}\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=v\}-P(w,v)\right)+\right.$
$\quad\left.\sum_{1\leq i^{\prime}\neq\ell^{\prime}\leq k}\mathbbm{1}\{X_{j-1}^{(i^{\prime})}=w\}\left(\mathbbm{1}\{X_{j}^{(\ell^{\prime})}=v\}-P(X_{j-1}^{(\ell^{\prime})},v)\right)\right)$
$=\sum_{w\in V}\sum_{i,i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}\left(\mathbbm{1}\{X_{j}^{(i)}=u\}-P(w,u)\right)$
$\qquad\qquad\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=v\}-P(w,v)\right)$
$+\sum_{w\in V}\sum_{i=1}^{k}\sum_{i^{\prime}\neq\ell^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}\cdot$
$\left(\mathbbm{1}\{X_{j}^{(i)}=u\}-P(w,u)\right)\left(\mathbbm{1}\{X_{j}^{(\ell^{\prime})}=v\}-P(X_{j-1}^{(\ell^{\prime})},v)\right)$
$+\sum_{w\in V}\sum_{i\neq\ell}\sum_{i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}\cdot$
$\left(\mathbbm{1}\{X_{j}^{(\ell)}=u\}-P(X_{j-1}^{(\ell)},u)\right)\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=v\}-P(w,v)\right)$
$+\sum_{w\in V}\sum_{i\neq\ell}\sum_{i^{\prime}\neq\ell^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}$
$\quad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell)}=u\}-P(X_{j-1}^{(\ell)},u)\right)$
$\quad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell^{\prime})}=v\}-P(X_{j-1}^{(\ell^{\prime})},v)\right).$
We now study its expectation $\mathop{\mathbb{E}}_{#1}[Y_{j}^{T}Y_{j}]$. In particular, we split it in four summations as above. We start with the first. Recall $X_{j-1}^{(i)}$ and $X_{j-1}^{(i^{\prime})}$ are independent for $i\neq i^{\prime}$. Therefore,
$A_{j}(u,v)$ $\coloneqq\mathbb{E}_{j-1}\Bigl{[}\sum_{w\in V}\sum_{i,i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}$
$\left(\mathbbm{1}\{X_{j}^{(i)}=u\}-P(w,u)\right)\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=v\}-P(w,v)\right)\Bigr{]}$
$=\sum_{w\in V}\sum_{i=1}^{k}\mathbb{E}_{j-1}\Bigl{[}\mathbbm{1}\{X_{j-1}^{(i)}=w\}\left(\mathbbm{1}\{X_{j}^{(i)}=u\}-P(w,u)\right)$
$\qquad\left(\mathbbm{1}\{X_{j}^{(i)}=v\}-P(w,v)\right)\Bigr{]}$
$=\sum_{i=1}^{k}\sum_{w\in V}\mathbbm{1}\{X_{j-1}^{(i)}=w\}\Bigl{(}\mathbbm{1}\{u=v\}P(w,u)(1-P(w,u))$
$\qquad-\mathbbm{1}\{u\neq v\}P(w,u)P(w,v)\Bigr{)}.$
Analogously, we have that
$B_{j}(u,v)\coloneqq\mathbb{E}_{j-1}\Bigl{[}\sum_{w\in V}\sum_{i=1}^{k}\sum_{i^{\prime}\neq\ell^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}$
$\left(\mathbbm{1}\{X_{j}^{(i)}=u\}-P(w,u)\right)\left(\mathbbm{1}\{X_{j}^{(\ell^{\prime})}=v\}-P(X_{j-1}^{(\ell^{\prime})},v)\right)\Bigr{]}$
$=\sum_{w\in V}\sum_{i\neq i^{\prime}}\mathbb{E}_{j-1}\Bigl{[}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}$
$\quad\cdot\left(\mathbbm{1}\{X_{j}^{(i)}=u\}-P(w,u)\right)$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(i)}=v\}-P(w,v)\right)\Bigr{]}$
$=\sum_{w\in V}\sum_{i\neq i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}$
$\qquad\cdot\Bigl{(}\mathbbm{1}\{u=v\}P(w,u)(1-P(w,u))$
$\qquad-\mathbbm{1}\{u\neq v\}P(w,u)P(w,v)\Bigr{)}.$
For the third sum,
$C_{j}(u,v)\coloneqq\mathbb{E}_{j-1}\Bigl{[}\sum_{w\in V}\sum_{i\neq\ell}\sum_{i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell)}=u\}-P(X_{j-1}^{(\ell)},u)\right)$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=v\}-P(w,v)\right)\Bigr{]}$
$=\mathbb{E}_{j-1}\Bigl{[}\sum_{w\in V}\sum_{i\neq\ell}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(\ell)}=w\}$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell)}=u\}-P(w,u)\right)$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell)}=v\}-P(w,v)\right)\Bigr{]}$
$=\sum_{w\in V}\sum_{i\neq\ell}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(\ell)}=w\}$
$\qquad\cdot\Bigl{(}\mathbbm{1}\{u=v\}P(w,u)(1-P(w,u))$
$\qquad\quad-\mathbbm{1}\{u\neq v\}P(w,u)P(w,v)\Bigr{)}.$
Finally, for the fourth sum,
$D_{j}(u,v)\coloneqq\mathbb{E}_{j-1}\Bigl{[}\sum_{w\in V}\sum_{i\neq\ell}\sum_{i^{\prime}\neq\ell^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell)}=u\}-P(X_{j-1}^{(\ell)},u)\right)$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell^{\prime})}=v\}-P(X_{j-1}^{(\ell^{\prime})},v)\right)\Bigr{]}$
$=\mathbb{E}_{j-1}\Bigl{[}\sum_{w\in V}\sum_{i,i^{\prime}\neq\ell}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell)}=u\}-P(X_{j-1}^{(\ell)},u)\right)$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell)}=v\}-P(X_{j-1}^{(\ell)},v)\right)\Bigr{]}$
$=\sum_{w,w^{\prime}\in V}\sum_{i,i^{\prime}\neq\ell}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w,X_{j-1}^{(\ell)}=w^{\prime}\}$
$\qquad\cdot\Bigl{(}\mathbbm{1}\{u=v\}P(w^{\prime},u)(1-P(w^{\prime},u))$
$\qquad\quad-\mathbbm{1}\{u\neq v\}P(w^{\prime},u)P(w^{\prime},v)\Bigr{)}$
$=\sum_{w\in V}\sum_{i,i^{\prime}\neq\ell}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})},X_{j-1}^{(\ell)}=w\}$
$\qquad\cdot\Bigl{(}\mathbbm{1}\{u=v\}P(w,u)(1-P(w,u))$
$\qquad\quad-\mathbbm{1}\{u\neq v\}P(w,u)P(w,v)\Bigr{)}.$
Notice that we have $W^{(1)}_{T}=\sum_{t=1}^{T}(A_{t}+B_{t}+C_{t}+D_{t})$. Moreover,
$\sum_{t=1}^{T}A_{t}(u,v)$ $=\mathbbm{1}\{u=v\}\sum_{w\in V}N(w)P(w,u)(1-P(w,u))$
$\quad-\mathbbm{1}\{u\neq v\}\sum_{w\in V}N(w)P(w,u)P(w,v).$
Moreover, by Lemma 3, with high probability, we have that $N(u)\leq 2kT\pi(u)$ for any $u\in V$. Therefore,
$\left\|\sum_{j=1}^{T}A_{t}\right\|_{1}$ $\leq 2\max_{u\in V}\sum_{w\in V}N(w)P(w,u)$
$\leq 4kT\max_{u\in V}\sum_{w\in V}\pi(w)P(w,u)$
$=4kT\pi^{\star},$
where the last equality follows from the fact that $\pi$ is the stationary measure. Essentially the same calculations show that
$\left\|\sum_{j=1}^{T}A_{t}\right\|_{\infty}\leq 4kT\pi^{\star},$
which implies that
$\left\|\sum_{j=1}^{T}A_{t}\right\|\leq\sqrt{\left\|\sum_{j=1}^{T}A_{t}\right\|_{1}\left\|\sum_{j=1}^{T}A_{t}\right\|_{\infty}}\leq 4kT\pi^{\star}$
To bound the contribution to the operator norm given by $B_{t}$, observe that by Lemma 3, w.h.p. and for any $u\in V$, $\widetilde{N}(u,u)\leq 2k(k-1)T\pi(u)$. Therefore, similarly as above, we have that
$\left\|\sum_{t=1}^{T}B_{t}\right\|_{1}$ $\leq 2\max_{u\in V}\sum_{w\in V}\widetilde{N}(w,w)P(w,u)$
$\leq 4k(k-1)T\max_{u\in V}\sum_{w\in V}\pi(w)P(w,u)$
$\leq 4k(k-1)T\pi^{\star}$
and
$\left\|\sum_{t=1}^{T}B_{t}\right\|_{\infty}\leq 4k(k-1)T\pi^{\star},$
which implies that
$\left\|\sum_{t=1}^{T}B_{t}\right\|\leq 4k(k-1)T\pi^{\star}.$
Analogously, we can show that
$\left\|\sum_{t=1}^{T}C_{t}\right\|\leq 4k(k-1)T\pi^{\star}.$
In a similar way we bound the contribution given by $D_{t}$. With high probability, by Lemma 3 it holds that,
$\left\|\sum_{t=1}^{T}D_{t}\right\|$ $\leq 2k^{2}\max_{u\in V}\sum_{w\in V}N(w)P(w,u)$
$\leq 4k^{3}T\pi^{\star}.$
By summing the contribution of $A_{t},B_{t},C_{t},D_{t}$ we prove the first part of the lemma.
We can obtain an analogous bound for $\|W^{(2)}_{t}\|$. We start by looking at each individual entry $Y_{j}Y_{j}^{T}(u,v)$
$Y_{j}Y_{j}^{T}(u,v)=\sum_{w\in V}Y_{j}(u,w)Y_{j}(v,w)$
$=\sum_{w\in V}\Bigl{(}\sum_{i=1}^{k}\mathbbm{1}\{X_{j-1}^{(i)}=u\}\left(\mathbbm{1}\{X_{j}^{(i)}=w\}-P(u,w)\right)$
$+\sum_{1\leq i\neq\ell\leq k}\mathbbm{1}\{X_{j-1}^{(i)}=u\}\left(\mathbbm{1}\{X_{j}^{(\ell)}=w\}-P(X_{j-1}^{(\ell)},w)\right)\Bigr{)}$
$\cdot\Bigl{(}\sum_{i^{\prime}=1}^{k}\mathbbm{1}\{X_{j-1}^{(i^{\prime})}=v\}\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=w\}-P(v,w)\right)$
$+\sum_{1\leq i^{\prime}\neq\ell^{\prime}\leq k}\mathbbm{1}\{X_{j-1}^{(i^{\prime})}=v\}\left(\mathbbm{1}\{X_{j}^{(\ell^{\prime})}=w\}-P(X_{j-1}^{(\ell^{\prime})},w)\right)\Bigr{)}$
$=\sum_{w\in V}\sum_{i,i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(i)}=w\}-P(u,w)\right)$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=w\}-P(v,w)\right)$
$\quad+\sum_{w\in V}\sum_{i=1}^{k}\sum_{i^{\prime}\neq\ell^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(i)}=w\}-P(u,w)\right)$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell^{\prime})}=w\}-P(X_{j-1}^{(\ell^{\prime})},w)\right)$
$\quad+\sum_{w\in V}\sum_{i\neq\ell}\sum_{i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(\ell)}=w\}-P(X_{j-1}^{(\ell)},w)\right)$
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=w\}-P(v,w)\right)$
$\quad+\sum_{w\in V}\sum_{i\neq\ell}\sum_{i^{\prime}\neq\ell^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\qquad\left(\mathbbm{1}\{X_{j}^{(\ell)}=w\}-P(X_{j-1}^{(\ell)},w)\right)$
$\qquad\left(\mathbbm{1}\{X_{j}^{(\ell^{\prime})}=w\}-P(X_{j-1}^{(\ell^{\prime})},w)\right).$
We now look at its expectation $\mathop{\mathbb{E}}_{#1}[Y_{j}^{T}Y_{j}]$. As done previously, we split it in four summations. By using the independence of $X_{j}^{(i)}$ from $X_{j}^{(i^{\prime})}$ for $i\neq i^{\prime}$, we have that
$A_{j}^{\prime}(u,v)$ $\coloneqq\mathbb{E}_{j-1}\Bigl{[}\sum_{w\in V}\sum_{i,i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\quad\left(\mathbbm{1}\{X_{j}^{(i)}=w\}-P(u,w)\right)\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=w\}-P(v,w)\right)\Bigr{]}$
$=\mathbbm{1}\{u=v,X_{j-1}^{(i)}=u\}\sum_{w\in V}\Bigl{(}P(u,w)(1-P(u,w))^{2}$
$\qquad+P(u,w)^{2}(1-P(u,w))\Bigr{)}$
$=\mathbbm{1}\{u=v,X_{j-1}^{(i)}=u\}\sum_{w\in V}P(u,w)(1-P(u,w)).$
Therefore,
$\sum_{t=1}^{T}A_{t}^{\prime}(u,v)$ $=\sum_{j}\mathbbm{1}\{u=v,X_{j-1}^{(i)}=u\}$
$\qquad\cdot\sum_{w\in V}\mathbbm{1}\{X_{j-1}^{(i)}=u\}P(u,w)(1-P(u,w))$
$\leq\mathbbm{1}\{u=v\}N(u)\sum_{w\in V}P(u,w)$
$=\mathbbm{1}\{u=v\}N(u).$
Hence,
$\left\|\sum_{t}A_{t}^{\prime}\right\|\leq\max_{u}N(u)\leq 2kT\pi^{\star}.$
Similarly,
$B_{j}^{\prime}(u,v)\coloneqq\mathbb{E}_{j-1}\Bigl{[}\sum_{w\in V}\sum_{i=1}^{k}\sum_{i^{\prime}\neq\ell^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\left(\mathbbm{1}\{X_{j}^{(i)}=w\}-P(u,w)\right)\left(\mathbbm{1}\{X_{j}^{(\ell^{\prime})}=w\}-P(X_{j-1}^{(\ell^{\prime})},w)\right)\Bigr{]}$
$=\sum_{w\in V}\sum_{i\neq i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\qquad\left(\mathbbm{1}\{X_{j}^{(i)}=w\}-P(u,w)\right)^{2}.$
Therefore,
$\left\|\sum_{t}B_{t}^{\prime}\right\|\leq\left\|\widetilde{N}\right\|\leq 2k(k-1)T\pi^{\star}.$
Similarly,
$C_{j}^{\prime}(u,v)\coloneqq\mathbb{E}_{j-1}\Bigl{[}\sum_{w\in V}\sum_{i\neq\ell}\sum_{i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\left(\mathbbm{1}\{X_{j}^{(\ell)}=w\}-P(X_{j-1}^{(\ell)},w)\right)\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=w\}-P(v,w)\right)\Bigr{]}$
$=\sum_{w\in V}\sum_{i\neq i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\qquad\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=w\}-P(v,w)\right)^{2},$
and
$\left\|\sum_{t}C_{t}^{\prime}\right\|\leq 2k(k-1)T\pi^{\star}.$
Finally,
$D_{j}^{\prime}(u,v)\coloneqq\sum_{w\in V}\sum_{i,i^{\prime}\neq\ell}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}\cdot$
$\qquad\mathbb{E}_{j-1}\Bigl{[}\left(\mathbbm{1}\{X_{j}^{(\ell)}=w\}-P(X_{j-1}^{(\ell)},w)\right)^{2}\Bigr{]}$
$=\sum_{w\in V}\sum_{i,i^{\prime}\neq\ell}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\qquad\Bigl{(}P(X_{j-1}^{(\ell)},w)(1-P(X_{j-1}^{(\ell)},w))^{2}$
$\qquad\quad+(1-P(X_{j-1}^{(\ell)},w))P(X_{j-1}^{(\ell)},w)^{2}\Bigr{)}$
$=\sum_{i,i^{\prime}\neq\ell}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
$\qquad\cdot\sum_{w\in V}P(X_{j-1}^{(\ell)},w)(1-P(X_{j-1}^{(\ell)},w))$
$\leq\sum_{i,i^{\prime}\neq\ell}\mathbbm{1}\{X_{j-1}^{(i)}=u,X_{j-1}^{(i^{\prime})}=v\}$
Therefore, by Lemma 3,
$\left\|\sum_{t=1}^{T}D_{t}^{\prime}\right\|\leq\left\|Q\right\|\leq 3Tk^{2}\pi^{\star}.$
By summing together these four contributions, we prove the second part of the lemma. ∎
We can now apply Theorem 4 with $R=\sqrt{2}k^{2}$ and $\sigma^{2}=8k^{3}T\pi^{\star}$. We obtain that
$\mathop{\mathbb{P}}[\exists k\geq 0\colon\|Y_{k}\|\geq r]\leq 2n\exp\left(-\frac{r^{2}/2}{8k^{3}T\pi^{\star}+\sqrt{2}k^{2}r/3}\right).$
In particular, let $r=(1/\sqrt{2})\cdot Tk\pi_{\star}\epsilon$ and choose $T\geq 5k^{3}\epsilon^{-2}\pi_{\star}^{-1}\log(n)(4+\|\pi\|_{2}^{2}\cdot\frac{\pi^{\star}}{\pi_{\star}})$, this failure probability will be $O(n^{-5})$.
We have proved that $\|N-(D_{N}+\widetilde{N})P\|=\|D_{N}(M-P-\widetilde{M}P))\|\leq(1/\sqrt{2})\cdot Tk\pi_{\star}\epsilon$ with high probability. By (5), we have that $\|D_{N}(M-P-(k-1)\Pi)\|\leq(5/6)\cdot Tk\pi_{\star}\epsilon$ with high probability. Since $\min_{u}D_{N}(u,u)\geq(9/10)\cdot Tk\pi_{\star}$ (again, with high probability), we have that $\|M-(P+(k-1)\Pi)\|\leq\epsilon$.
8 Acknowledgements
All authors were partially supported by the Defence Science and Technology Laboratory.
[…]

### v2

@@
-Wireless Network Topology Inference: A Markov Chains Approach
+
 Abstract
-In this work, we address the problem of inferring the topology of a wireless network using limited observational data.
+We address the problem of inferring the topology of a wireless network using limited observational data.
 Specifically, we assume that we can detect when a node is transmitting, but no further information regarding the transmission is available.
@@
 Applying this estimation procedure to wireless networks, our numerical experiments demonstrate that the proposed method accurately infers network topology across a wide range of parameters, consistently outperforming transfer entropy, particularly under conditions of high network congestion.
+{IEEEkeywords}
+Markov chains, topology inference, wireless networks.
 1 INTRODUCTION
-Inferring the topology of a network from limited information has numerous applications across various fields, including the discovery of gene interactions in computational biology, the analysis of climate dynamics in environmental science, and the investigation of functional brain structure in neuroscience [net_inf_survey].
+Inferring the topology of a network from limited information has numerous applications across various fields, including the discovery of gene interactions in computational biology, the analysis of climate dynamics in environmental science, and the investigation of functional brain structure in neuroscience [1, 2, 3].
 Additionally, in defense and security contexts, understanding the structure of an adversary’s network can provide valuable insights for decision-making.
@@
 The objective is then to infer whether pairs of nodes are connected by exploiting correlations in their corresponding time series.
-Previous work, particularly in the context of wireless networks [TR13, LC17, TG20, liu2022topology], has focused on detecting links by computing causality measures between time series, such as Granger causality [granger69] and transfer entropy [schreiber00].
+Previous work, particularly in the context of wireless networks [4, 5, 6, 7], has focused on detecting links by computing causality measures between time series, such as Granger causality [8] and transfer entropy [9].
 While these approaches have demonstrated reasonable accuracy, particularly in simulations, they remain poorly understood.
@@
 Specifically, we monitor a network of wireless devices, assuming we can detect when a device is transmitting, but not the target of the transmission.
-Such data could be collected via a network of radio frequency sensors.
+Such data could be collected via a network of radio frequency sensors (see Fig.
+1 for a schematic depiction).
+Figure 1: A network of wireless devices (blue circles) monitored by a collection of radio frequency sensors (green squares).
+Whenever a wireless device is transmitting, nearby sensors will measure an associated power bursts.
+Our goal is to infer the topology of the network using only the sensors’ measurements.
 To develop our algorithm, we adopt a Markovian assumption regarding the network’s dynamics: we assume that the future state of the network depends only on its current state and is independent of its past states.
@@
 Our primary theoretical contribution is the development of a consistent estimator.
-We prove that by observing the trajectories of the chains for a sufficient duration, we can construct an estimated transition matrix $\hat{P}$ that approximates the true matrix $P$ arbitrarily well in the operator norm.
-Crucially, this approximation in operator norm provides insights into the global structure of the network, such as the presence of clusters and bottlenecks [STsparse].
-We apply this estimation procedure to wireless networks simulated using the ns-3 network simulator [ns3].
+We prove that, by observing the trajectories of the chains for a sufficient duration, we can construct an estimated transition matrix $\hat{P}$ that approximates the true matrix $P$ arbitrarily well in the operator norm.
+Crucially, this approximation in operator norm provides insights into the global structure of the network, such as the presence of clusters and bottlenecks [10].
+We apply this estimation procedure to wireless networks simulated using the ns-3 network simulator [11].
 Our experimental results show that the proposed method is competitive with state-of-the-art techniques.
@@
 1.2 Related Work
-From a theoretical standpoint, our work relates to the broader problem of estimating Markov chain parameters from observed trajectories, an area that has garnered significant attention in recent years [hsu19, Wolf19, WolfKont24].
-Our algorithm builds upon techniques developed by Wolfer and Kontorovich [Wolf19].
-However, prior work typically focuses on scenarios where only a single Markov chain is observed, which simplifies the estimation task but is less applicable to modelling topology inference in wireless networks.
+From a theoretical standpoint, our work relates to the broader problem of estimating Markov chain parameters from observed trajectories, an area that has garnered significant attention in recent years [12, 13, 14].
+Our algorithm builds upon techniques developed by Wolfer and Kontorovich [13].
+Prior work, however, typically focuses on scenarios where only a single Markov chain is observed, which simplifies the estimation task but is less applicable to modelling topology inference in wireless networks.
 In the single-chain case, observing consecutive transmissions at nodes $u$ and $v$ directly implies the existence of a link between them, a property that does not hold in our model.
-There is a significant body of literature on wireless network topology inference (see, e.g.,  [TR13, LC17, salmond19, TG20, mehrotra2021minimax, liu2022topology, du2023network]).
-However, to the best of our knowledge, our work is the first to provide theoretical guarantees for the estimation procedure, albeit under a simplified model of network dynamics.
+There is a significant body of literature on wireless network topology inference (see, e.g.,  [4, 5, 15, 6, 16, 7, 17]).
+However, to the best of our knowledge, our work is the first to provide theoretical guarantees for the estimation procedure, at least under a simplified model of network dynamics.
 While this model may not fully capture the complexities of real wireless networks, we believe it encapsulates the key challenges inherent to the inference task.
@@
 We discuss potential directions for future research in the conclusion section.
+1.3 Organisation
+We introduce the setup and the formal model underpinning our inference procedure in Section 2, which then is described in Section 3.
+In Section 4 we state and discuss our main theoretical result, while its proof is deferred to Section 7.
+We showcase and discuss our experimental results in Section 5.
+In particular, we first consider an idealised scenario in which we know exactly when a given node is transmitting (Section 5 5.1) and we later discuss how to acquire this information using radio frequency sensors (Section 5 5.2).
+We end with a discussion of possible future work in Section 6.
+Code and datasets to reproduce our experiments are available at [18].
 2 PROBLEM SETUP
@@
 Our main task is to infer the set of edges $E$ given the time series matrix $TS$.
+We remark that, in practical scenarios, we do not usually have access to perfect information about when a device is transmitting.
+Rather, we can only obtain an approximation of the time series matrix, for example, by deploying a set of radio frequency sensors in the environment.
+We expand on how RF sensors can be used to approximate this time series matrix in Section 5 5.2.
 To formally model network dynamics, we consider $k$ independent Markov chains $X^{(1)},X^{(2)},\dots,X^{(k)}$ on a finite state space $V$ (the set of nodes of our network) and with the same transition matrix $P$.
@@
 For technical reasons, we assume $P$ has a unique stationary distribution $\pi$, i.e., $\pi P=\pi$.
-This is a reasonable assumption in our context since it essentially means the network is (strongly) connected.
-Furthermore, we assume the Markov chains are stationary, i.e., $X^{(i)}_{0}\sim\pi$ for all $1\leq i\leq k$.
+This is a reasonable assumption in our context since it essentially means the network is (strongly) connected, i.e., any two nodes of the network can communicate with each other (without necessarily being connected by a link).
+Furthermore, we assume the Markov chains are stationary, i.e., $X^{(i)}_{1}\sim\pi$ for all $1\leq i\leq k$.
 This technical assumption is made to simplify our analysis, but is not necessary and could be dropped with additional technical work (without modifying the proposed algorithm).
@@
 Indeed, our goal is to estimate $P$.
-What we are able to observe is, for each time step $t=0,\dots,T$, how many chains are currently visiting a given vertex $v\in V$.
-In other words, for any $t=1,\dots,T$ and $v\in V$, we can observe the quantity $S_{t,v}=|\{i\colon X^{(i)}_{t}=v\}|$ for any $t\in\mathbb{N}_{+},v\in V$.
+What we are able to observe is, for each time step $t=1,\dots,T$, how many chains are currently visiting a given vertex $v\in V$.
+In other words, for any $t=1,\dots,T$ and $v\in V$, we can observe the quantity $S_{t,v}=|\{i\colon X^{(i)}_{t}=v\}|$.
 We are not able, however, to observe if a specific chain is visiting $v$ at time $t$.
@@
 $N(u,v)$ =|$=\left|\{i,j\in[k],t\in[T]\colon X^{(i)}_{t-1}=u,X^{(j)}_{t}=v\}\right|,$
-$N(u)$ =|$=\left|\{i\in[k],0\leq t<T\colon X^{(i)}_{t}=u\}\right|.$
+$N(u)$ =|$=\left|\{i\in[k],1\leq t<T\colon X^{(i)}_{t}=u\}\right|.$
 In other words, $N(u,v)$ represents the number of times we have observed $u$’s transmission being immediately followed by $v$’s, while $N(u)$ counts the total number of times $u$ is transmitting.
 Notice that $N(u)=\sum_{v\in V}N(u,v)$.
-Moreover, these quantities can be computed having access only to $S_{t,v}$ for $0\leq t\leq T,v\in V$.
+Moreover, these quantities can be computed having access only to $S_{t,v}$ for $1\leq t\leq T,v\in V$.
 Let also $M\in\mathbb{R}^{n\times n}$ be defined as
@@
 $\hat{P}=M-(k-1)\widehat{\Pi}.$
-The term $(k-1)\widehat{\Pi}$ is a correction necessary because nodes that are visited by the chains more frequently will have larger entries in $M$.
-Notice that computing $\hat{P}$ is computationally efficient and can be done in an online manner: we do not have to store the entire time series matrix $TS$, we can just update $N(u,v)$, $N(v)$, and $\hat{P}$ at each time step.
+The term $(k-1)\widehat{\Pi}$ is a correction necessary because nodes that are visited by the chains more frequently will have larger entries in $M$; in particular, if the chains often visit two nodes $u$ and $v$, the corresponding entry $M(u,v)$ will be large even though there might not be an edge between $u$ and $v$.
+Indeed, we expect that, on average, $M(u,v)\approx(k-1)\widehat{\Pi}$ whenever $u$ is not connected to $v$.
+Computing $\hat{P}$ is computationally efficient and can be done in an online manner: we do not have to store the entire time series matrix $TS$, we can just update $N(u,v)$, $N(v)$, and $\hat{P}$ at each time step.
+This can be done in time $O(n^{2}T)$ in total.
 Returning to the wireless network setting, $\hat{P}(u,v)$ can be interpreted as our estimated probability that $v$ will transmit immediately after $u$ has transmitted.
 While $\hat{P}(u,v)$ is strongly correlated to the likelihood $u$ and $v$ are connected by a link, it is not a measure of how likely a link between $u$ and $v$ exists.
-To this end, we suggest using an estimator for the Laplacian associated to $P$.
+In particular, if a node $u$ has many outgoing links, entries in the corresponding row $\hat{P}(u,\cdot)$ will necessarily be small.
+To overcome this issue, we suggest using an estimator for the Laplacian associated to $P$.
 For a transition matrix $P$ with stationary distribution $\pi$, the Laplacian $L$ of $P$ is the matrix with entries $L(u,v)=P(u,v)\sqrt{\pi(v)/\pi(u)}$.
-We recall that any Markov chain can be interpreted as a random walk on a weighted directed graph, where $L(u,v)$ represents the weight on the edge $(u,v)$ [markovmixing].
+We recall that any Markov chain can be interpreted as a random walk on a weighted directed graph, where $L(u,v)$ represents the weight on the edge $(u,v)$ [19].
 Therefore, to infer if a link is present or not, we will use the estimated matrix $\hat{L}$, defined as
@@
 where $\hat{\pi}$ is the leading eigenvector of $\hat{P}$ and an estimate for the stationary distribution $\pi$.
-Finally, when we know that links in the observed network are bidirectional, it is natural to consider a symmetrisation of the Laplacian instead:
+When we know that links in the observed network are bidirectional, it is natural to consider a symmetrisation of the Laplacian [20] instead:
 $\hat{L}_{\text{sym}}(u,v)=\sqrt{\frac{\hat{\pi}(v)}{\hat{\pi}(u)}}\cdot\frac{\hat{P}(u,v)+\hat{P}(v,u)}{2}.$
-We conclude this section by underlining that our estimator only records transmissions happening in two consecutive time intervals.
-This means it’s important to choose an appropriate interval length, otherwise network delays might severely impact our estimator.
+We conclude this section by underlining two issues that arise when applying our methodology to infer the topology of real-world wireless networks.
+First of all, notice that our estimator only records transmissions happening in two consecutive time intervals.
+This means it is important to choose an appropriate interval length, otherwise network delays might severely impact our estimator.
 In our simulations, however, this constraint has not been detrimental to our results.
 The situation might be more complicated in real world settings.
-We plan to study how to bypass these potential issues in future work.
+Secondly, our estimator depends on the number $k$ of Markov chains that are concurrently active.
+In practice, $k$ can be interpreted as the average number of devices that transmit in the same time interval.
+We will estimate this parameter as the ratio between the number of nonzero entries in the time series matrix $TS$ and the number of time-intervals $t$ such that there has been a transmission both at interval $t$ and at the following interval $t+1$ (this is because our estimator $\hat{P}$ only considers consecutive transmissions).
 4 THEORETICAL GUARANTEES
@@
 This approximation preserves important quantities of the network.
-For example, it approximately preserves its stationary distribution; this tells us how often, on average, nodes are visited [markovmixing], which corresponds how often a node is transmitting in our model.
-It also preserves its spectral gap, which tells us how fast a Markov chain with transition matrix $P$ converges to its stationary distribution from a worst case starting point [MonteTetali].
-Moreover, it preserves important connectivity properties of the networks, such as its “cut-structure”, that is, the structure of clusters and bottlenecks in the network [PSZ, STsparse].
+For example, it approximately preserves its stationary distribution; this tells us how often, on average, nodes are visited [19], which corresponds how often a node is transmitting in our model.
+It also preserves its spectral gap, which tells us how fast a Markov chain with transition matrix $P$ converges to its stationary distribution from a worst case starting point [20].
+Moreover, it preserves important connectivity properties of the networks, such as its “cut-structure”, that is, the structure of clusters and bottlenecks in the network [21, 10].
 What this approximation doesn’t necessarily preserve are the individual links in the network.
 For example, consider a transition matrix of a random walk on a complete graph of $n$ vertices.
-Now consider an Erdos-Renyi random graph $G(n,p)$: that is, a graph of $n$ vertices where we connect any two vertices with probability $p$.
-As long as $p\gg log(n)/n$, it is well-known that the transition matrix of a random walk on $G(n,p)$ well approximates (in the operator norm) the transition matrix of the complete graph [STsparse].
+Now consider an Erdős-Rényi random graph $G(n,p)$: that is, a graph of $n$ vertices where we connect any two vertices with probability $p$.
+As long as $p\gg log(n)/n$, it is well-known that the transition matrix of a random walk on $G(n,p)$ well approximates (in the operator norm) the transition matrix of the complete graph [10].
 However, the $G(n,p)$ graph might have only $O(n\log(n))$ edges compared to the $\Theta(n^{2})$ edges of the complete graph, i.e., only a small fraction of the edges are preserved.
@@
 In particular, it is inversely linear in the spectral gap $\lambda(P)$.
-The spectral gap is a crucial parameter in the analysis of Markov chains: its reciprocal, $1/\lambda(P)$ provides a bound on the mixing time of $P$, i.e., the time it takes for a Markov chain with transition matrix $P$ to converge to the stationary distribution from a worst case starting point [MonteTetali].
-A long as $P$ is irreducible, as in our case, $\lambda(P)>0$.
+The spectral gap is a crucial parameter in the analysis of Markov chains: its reciprocal, $1/\lambda(P)$ provides a bound on the mixing time of $P$, i.e., the time it takes for a Markov chain with transition matrix $P$ to converge to the stationary distribution from a worst case starting point [20].
+As long as $P$ is irreducible, as in our case, $\lambda(P)>0$.
 We believe that, in cases where $P$ models a communication network, $1/\lambda(P)$ is relatively small.
 Furthermore, for natural transition matrices $P$, $1/\pi_{\star}$ is proportional to $n$, the size of the network.
-A larger value of $1/\pi_{\star}$ would mean, in our model, that there are nodes that transmits very rarely: to approximate $P$, we necessarily need to collect enough information about these nodes, increasing the time required.
+A larger value of $1/\pi_{\star}$ would mean, in our model, that there are nodes that transmit very rarely: to approximate $P$, we necessarily need to collect enough information about these nodes, increasing the time required.
 Finally, $\frac{\pi^{\star}}{\pi_{\star}}$ can be interpreted as the ratio between the largest and the smallest frequencies of transmission of a node: if nodes transmit at roughly the same frequency, this parameter becomes order of 1.
 In summary, for many network topologies, as long as the number of concurrent transmissions $k$ is not too large, Theorem 1 asserts that a number of samples proportional to the size of the network suffices to well approximate $P$.
-The proof of Theorem 1, deferred to Section 7, is inspired by the techniques of [Wolf19].
-In particular, it uses matrix concentration inequalities for Markov chains [matrixMarkovBernstein] and martingales [Tropp].
+The proof of Theorem 1, deferred to Section 7, is inspired by the techniques of [13]. In particular, it uses matrix concentration inequalities for Markov chains [22] and martingales [23].
 We end this section by briefly discussing the validity of our Markovian model for capturing our task of estimating the topology of a wireless network.
@@
 For this reason, our Markovian model can only be realistic if the chance of two or more chains visiting the same node at the same time is very small.
-Formally, the probability that at a time $t$ two stationary chains visit the same node can be upper bounded by
+Formally, the probability that at some time $t$ two stationary chains visit the same node can be upper bounded by
 $\sum_{u\in V}\sum_{1\leq i<j\leq k}\pi(u)={k\choose 2}\sum_{u\in V}\pi(u)^{2}\leq k^{2}\|\pi\|_{2}^{2}.$
@@
 5 EXPERIMENTAL RESULTS
-We have simulated various wireless network topologies with the network simulator ns-3 [ns3].
+We have simulated various wireless network topologies with the network simulator ns-3 [11].
 Each wireless network is constructed by generating $n$ points on the plane.
@@
 Different point sets will induce a different network topology.
-Propagation speed is constant (the speed of light) and path loss is dependent only on the transmission range: if the receiver is within range, it will receive the transmission at 16dbm, else it will not receive a transmission.
 We use the IEEE 802.11b protocol with a data rate of 1Mbps.
@@
 More precisely, we will consider the matrix $M_{\text{TE}}$ where the $(i,j)$-entry is equal to the transfer entropy from the time series associated with $i$ to the time series associated with $j$.
-To make the comparison even, we will further symmetrised this matrix by taking the average between this matrix and its transpose.
+To make the comparison even, we will also symmetrise this matrix by taking the average between this matrix and its transpose.
+Transfer entropy is estimated using the Python library PyInform [24] with embedding dimension $d=5$.
+We did not optimise such number but found a similar performance for smaller values of $d$, while larger values of $d$ might produce worse results.
+5.1 An idealised setting
+We first consider an idealised setting that follows the problem setup described in Section 2: we assume we have direct access to the time series matrix $TS$.
+While we do not expect to have perfect information about the time series matrix $TS$ in practice, we believe useful to know the strengths and limitations of our approach in this idealised scenario.
+We fix a constant propagation speed (the speed of light) and path loss is dependent only on the transmission range: if the receiver is within range, it will receive the transmission at 16dbm, else it will not receive a transmission.
 In Fig.
@@
 In Fig.
-1(d) we stud the effect of a varying transmission length on network estimation.
+2(d) we study the effect of a varying transmission length on network estimation.
 In particular, we record the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window.
@@
 transfer entropy) if it appears in the top $m$ entries of the upper triangular part of the $\hat{L}_{\text{sym}}$ (resp.
-transfer entropy) matrix, where $m$ in the number of links in the network.
+transfer entropy) matrix, where $m$ is the number of links in the network.
+[^1] We have adopted this simplistic inference procedure only to compare our estimator to transfer entropy.
+More complicated inference procedures, such as ones based on surrogate data approaches [25] or machine learning methods [6] might be needed in real world scenarios.
 We observe that a shorter transmission window makes estimation harder since it implies a larger number of nodes will be concurrently transmitting (i.e., a larger parameter $k$ in 1).
@@
 Further experimental results are reported in Figs.
-2, 3, 4 and 4.
-We highlight our methodology consistently outperforms transfer entropy whenever the network is congested.
-Transfer entropy and our estimate appear very similar when $T\geq 15s$.
+3, 4, 5 and 6.
+We highlight that our methodology consistently outperforms transfer entropy whenever the network is congested.
+Transfer entropy and our estimate appear to work similarly well when $T\geq 15s$, with our estimator, however, giving more weight to pairs of nodes not connected by links: we believe this might be due to an underestimation of the value of $k$ defined in Section 3.
+A more refined approach for estimating $k$ might further improve our estimation procedure.
 Figure 2: Experimental results for a cyclic network topology of six nodes.
@@
 All results are averaged over $50$ independent simulations.
-Figure 2: Experimental results for a network topology of eight nodes.
-Figure 3: Experimental results for a network topology where nodes are arranged in a three-by-three grid.
-Figure 4: Experimental results for a network topology where nodes are arranged in a four-by-two grid.
-Figure 5: Experimental results for a network topology of twenty nodes arranged in a path.
+A description of the experimental setup is provided at the beginning of Section 5.
+Figure 3: Experimental results for a network topology of eight nodes. See Fig. 2 for a description.
+Figure 4: Experimental results for a network topology where nodes are arranged in a three-by-three grid. See Fig. 2 for a description.
+Figure 5: Experimental results for a network topology where nodes are arranged in a four-by-two grid. See Fig. 2 for a description.
+Figure 6: Experimental results for a network topology of twenty nodes arranged in a path. See Fig. 2 for a description.
+5.2 A more realistic setting
+We now consider a more realistic scenario in which we do not have perfect information about the time series matrix TS.
+For simplicity, we will assume to know the number of nodes in the network, but not their position.
+To obtain information about packet transmission, we assume we are able to deploy a set of $s$ RF sensors in the environment.
+When a packet is transmitted from a node $u$, RF sensors positioned close to $u$ will be able to detect the corresponding power burst.
+We use the Log Distance path loss model in our ns-3 simulations: this implies that, the farther away from $u$ a sensor $\sigma$ is, the more attenuated the received signal by the sensor.
+We again divide the time into small intervals $\{1,\dots,T^{\prime}\}$ (we assume a power burst is detected by at least a sensor at each interval, and discard intervals in which no power burst is detected); we can then construct a matrix $PD\in\mathbb{R}^{m\times T^{\prime}}$, where $PD(\sigma,t)$ represents the power detected at time $t$ by sensor $\sigma$.
+Suppose that, for a pair of time intervals $t,t^{\prime}$, the two vectors $PD(\cdot,t)$ and $PD(\cdot,t^{\prime})$ are somewhat similar: this indicates that the same node might be responsible for the detected power burst.
+This observation suggests a natural way to discern if the same node is transmitting during different time intervals: partition the vectors $\{PD(\cdot,t)\}_{t=1}^{T^{\prime}}$ using a clustering technique such as $k$-means.
+Time intervals assigned to the same cluster corresponds to power bursts we associate to the same node.
+In this way it is possible to construct a matrix which, hopefully, approximates the time series matrix TS (since we do not assume any information about the identity of the nodes, such as their position, this is only possible up to a relabelling of the nodes).
+We perform experiments using the same ns-3 experimental setup discussed at beginning of the section, but instead of fixing the network topology we place $n$ transmitters and $s$ RF sensors at random in a rectangular box.
+For example, in Fig.
+7 we have randomly placed $25$ transmitters and $100$ sensors in a square box of side $120m$.
+Only transmitters placed at most $50m$ apart are connected by a link in the network.
+This gives rise to a network topology with adjacency matrix displayed in Fig.
+7(d).
+We simulate the network behaviour with ns-3 exchanging $500$ messages between randomly chosen pair of nodes.
+As explained above, we use $k$-means clustering (with $k=n$) to group together power bursts measured by different sensors in the same interval.
+Since we do not have any information about the identity of the transmitters (recall, we do not assume any information about the location of the sensors or transmitters), to evaluate how well measured power bursts are clustered together, we label each cluster computed by $k$-means according to the transmitter that corresponds to the largest number of bursts in the cluster.
+We display in Fig.
+7(b) the percentage of power bursts detected that are associated to the correct transmitter for different lengths of the transmission window.
+The results are averaged over $20$ random realisations in which we fix the same network topology but randomly vary the position of the sensors.
+As expected, a larger transmission window results in a higher accuracy thanks to a lower number of concurrent transmissions.
+We then compute our estimator $\hat{L}_{\text{sym}}$ together with the transfer entropy matrix $M_{\text{TE}}$.
+Examples of these two matrices for a particular random realisation are shown in Fig.
+7 (e) and (f).
+Notice how these two matrices appear quite similar.
+In Fig.
+7(c) we compare the performances of our method with transfer entropy.
+Again, we measured the number of times the top $m$ entries in the two matrices correspond to links in the network, where $m$ is the number of links in the network.
+Both methodologies appear to perform reasonably well for $T\geq 10$, with our estimation procedure slightly but consistently improving upon transfer entropy.
+We repeat the same experiments with a bigger network: we randomly place $50$ transmitters and $150$ sensors in a rectangular box of $340m\times 200m$.
+We simulate the network with ns-3 exchanging now $2000$ messages between randomly chosen pairs.
+We perform $5$ random realisations, randomly placing both transmitters and sensors.
+Experimental results are shown in Fig.
+8.
+We can observe a degradation of the performances, both in the proportion of identified transmitters (see Fig.
+8(b)) and in the proportion of links of the network that corresponds to the top $m$ entries in both transfer entropy and $\hat{L}_{\text{sym}}$ matrix (see Fig.
+8(c)).
+The first observation suggests that, in large networks, a simple application of $k$-means might not be enough to identify the correct transmitter for most power bursts; future work might want to exploit potential information about the position of sensors or different source separation procedures such as using independent component analysis [6].
+The decrease in performance in correctly identifying links is not unexpected: the number of possible links increases quadratically in the number of transmitters while the number of actual links increases, roughly, linearly.
+Nevertheless, to help us better understand the reasons behind this decrease in performance, let us study closely Fig.
+8 (f), which compares the top entries of $\hat{L}_{\text{sym}}$ against the actual links in the network: first of all, notice how the main “backbone” of the network is essentially preserved.
+As suggested by our theoretical analysis, our method does not necessarily preserve individual links but rather structural properties of the network.
+Furthermore, false positive links typically connect nodes that are only two hops apart in the network.
+This is not surprising since if there exist links $\{u,v\}$ and $\{v,w\}$, the time series corresponding to $u$ and $w$ will be correlated.
+This suggests it might be possible to reduce this error by, for example, combining information about time series with different interval lengths.
+We leave this as future work.
+Figure 7: Experimental results for a random network constructed by placing $25$ transmitters and $100$ RF sensors at random in a $120m\times 120m$ square box.
+Figure (a) displays the position of the transmitters (black circles) and of the sensors (red crosses) for a particular random realisation, together with the resulting network topology.
+Figure (b) shows the number of power bursts correctly associated to the corresponding transmitters for a varying transmission window length (we display the average number over 20 random realisations, together with a 95% confidence interval).
+Figure (c) represents the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window (averaged over 20 random realisations).
+Figure (d) displays the adjacency matrix of the network.
+Figure (e) the estimated matrix $\hat{L}_{\text{sym}}$ for a transmission window length $T=10s$.
+Figure (f) the transfer entropy matrix ($T=10s$).
+Figure 8: Experimental results for a random network constructed by placing $50$ transmitters and $150$ RF sensors at random in a $340m\times 200m$ rectangular box.
+Figure (a) displays the adjacency matrix of the network for a random realisation.
+Figure (b) an estimated matrix $\hat{L}_{\text{sym}}$ for a transmission window length $T=60s$.
+Figure (c) the corresponding transfer entropy matrix ($T=60s$).
+Figure (d) shows the number of power bursts correctly associated to the corresponding transmitters for a varying transmission window length (we display the average number over 5 random realisations together with a 95% confidence interval).
+Figure (e) represents the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window (averaged over 5 random realisations).
+Figure (f) displays (for a specific random realisation with window length $T=60s$) the correctly identified links by $\hat{L}_{\text{sym}}$ in green, and false positives in red; their thickness is proportional to the corresponding entry of $\hat{L}_{\text{sym}}$.
+Dashed black lines correspond to links in the network that do not appear among the top entries of $\hat{L}_{\text{sym}}$.
 6 CONCLUSION AND FUTURE WORK
-In this work, we introduced a novel technique for estimating the topological structure of wireless networks.
-We demonstrated that our approach provably recovers information about network topology under a simplified Markovian model of network dynamics.
+In this work, we have introduced a novel technique for estimating the topological structure of wireless networks.
+We demonstrated that our approach provably recovers information about the network topology under a simplified Markovian model of network dynamics.
 Additionally, our experimental results showed that the proposed method is competitive with transfer entropy in realistic network simulations, consistently outperforming it under conditions of network congestion.
@@
 Finally, measures such as transfer entropy and Granger causality are often used as features in more sophisticated estimation procedures.
-For instance, Testi and Giorgetti [TG20] utilise these measures to train a neural network model.
+For instance, Testi and Giorgetti [6] utilise these measures to train a neural network model.
 It would be interesting to explore the integration of our techniques into machine learning frameworks for network inference.
 7 PROOFS
-We will show that, for $T$ large enough, $M(u,v)\approx P(u,v)-(k-1)\pi(v)$ and $N(v)\approx T\cdot k\cdot\pi(v)$.
+In this section we provide a proof of our main theoretical result, 1.
+Informally, we will show that, for $T$ large enough, $M(u,v)\approx P(u,v)-(k-1)\pi(v)$ and $N(v)\approx T\cdot k\cdot\pi(v)$.
 Therefore, by computing $M$ and $N$, we are able to approximate $P$.
@@
 For technical reason, we also need a bound on the following quantities ($u,v\in V$):
-$\widetilde{N}(u,v)=\sum_{t=0}^{T-1}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\},$ (2)
-$Q(u,v)=\sum_{t=0}^{T-1}\sum_{i,i^{\prime}}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(i^{\prime})}=v\}.$ (3)
-We will use the following recent matrix Bernstein inequality for Markov chains [matrixMarkovBernstein].
+$\widetilde{N}(u,v)=\sum_{t=1}^{T}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\},$ (2)
+$Q(u,v)=\sum_{t=1}^{T}\sum_{i,i^{\prime}}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(i^{\prime})}=v\}.$ (3)
+We will use the following recent matrix Bernstein inequality for Markov chains [22].
 Theorem 2.
@@
 $\mathop{\mathbb{P}}[\left\|\sum_{t=1}^{T}f(X_{t})\right\|>\epsilon]\leq 2d^{2}\exp\left(-\frac{\epsilon^{2}\lambda/72}{T\sigma^{2}+A\epsilon}\right).$
+We use 2 to bound the probability $N$, $\tilde{N}$, and $Q$ deviate from their expectation.
 Lemma 3.
@@
 Let $D_{N}$ be the $n\times n$ diagonal matrix such that $D_{N}(u,u)=N(u)$.
-Then, with probability at least $1-\delta$, for all $u,v\in V$, the following holds:
+Then, with probability at least $1-\delta$, for all $v\in V$, the following holds:
 $\left|N(v)-kT\pi(v)\right|\leq\epsilon kT\pi(v),$
 $\left\|\widetilde{N}-Tk(k-1)\pi\pi^{T}\right\|\leq\epsilon Tk(k-1)\pi_{\star}$
-$\left\|Q-Tk(k-1)\pi\pi^{T}-D_{N}\right\|\leq Tk^{2}\pi^{\star}$
+$\left\|Q-Tk(k-1)\pi\pi^{T}-D_{N}\right\|\leq Tk^{2}\pi^{\star}.$
 Proof.
 We first fix $v\in V$ and compute the expectation of $\mathop{\mathbb{E}}[N(v)]$:
-$\mathop{\mathbb{E}}[N(v)]$ $=\sum_{i=1}^{k}\sum_{t=0}^{T-1}\mathop{\mathbb{P}}[X^{(j)}_{t}=v]=kT\pi(v),$
+$\mathop{\mathbb{E}}[N(v)]$ $=\sum_{i=1}^{k}\sum_{t=1}^{T}\mathop{\mathbb{P}}[X^{(j)}_{t}=v]=kT\pi(v),$
 where the last equality holds because we assume the chains start from stationarity.
@@
 For the second inequality, we first compute $\mathop{\mathbb{E}}[\widetilde{N}(u,v)]$:
-$\mathop{\mathbb{E}}[\widetilde{N}(u,v)]$ $=\sum_{t=0}^{T-1}\sum_{i\neq j}\mathop{\mathbb{E}}[\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\}]$
+$\mathop{\mathbb{E}}[\widetilde{N}(u,v)]$ $=\sum_{t=1}^{T}\sum_{i\neq j}\mathop{\mathbb{E}}[\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\}]$
 $=k(k-1)T\pi(u)\pi(v).$
@@
 $g_{ij}((u,v))(x,y)=\begin{cases}1-\pi(u)\pi(v)&\text{ if }\{u,v\}=\{x,y\}\\ -\pi(x)\pi(y)&\text{ otherwise.}\end{cases}$
-Notice that $\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}]=\sum_{i\neq j}\sum_{t=0}^{T-1}g_{ij}(X_{t}^{(i)},X_{t}^{(j)})$.
+Notice that $\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}]=\sum_{i\neq j}\sum_{t=1}^{T}g_{ij}(X_{t}^{(i)},X_{t}^{(j)})$.
 Clearly, $\|g_{ij}((u,v))\|\leq 2$ for all $(u,v)\in V\times V$.
-Furthermore, simple calculations show that
-$\left\|\mathop{\mathbb{E}}_{#1}[g_{ij}((u,v))^{2}]\right\|\leq 2\pi^{\star}.$
-Therefore, with probability at least $\delta/k^{2}$, Theorem 2 implies that
-$\left\|\sum_{t=0}^{T-1}g_{ij}(X_{t}^{(i)},X_{t}^{(j)})\right\|\leq Tk\pi_{\star}.$
+Furthermore, simple calculations show that $\|\mathop{\mathbb{E}}_{#1}[g_{ij}((u,v))^{2}]\leq 2\pi^{\star}\|$.
+Therefore, with probability at least $\delta/k^{2}$, Theorem 2 implies that $\left\|\sum_{t=1}^{T}g_{ij}(X_{t}^{(i)},X_{t}^{(j)})\right\|\leq Tk\pi_{\star}$.
 A union bound over all $1\leq i\neq j\leq k$ yields the second statement.
-Finally, the third inequality follows simply from the first two by noticing that $Q(u,v)=\widetilde{N}(u,v)+D_{N}(u,v)$, ∎
+Finally, the third inequality follows simply from the first two by noticing that $Q(u,v)=\widetilde{N}(u,v)+D_{N}(u,v)$. ∎
 We now prove concentration for $M$.
-We apply similar techniques as in Kontorovich and Wolfer [Wolf19].
-We will need the following concentration inequality for matrix martingales, which is due to Tropp [Tropp].
+We apply similar techniques as in Kontorovich and Wolfer [13].
+We will need the following concentration inequality for matrix martingales, which is due to Tropp [23].
 Theorem 4.
@@
 $\mathop{\mathbb{E}}_{#1}[Y_{k}]=Y_{k-1}\quad\text{ for all }k=1,2,\dots,$
-where with $\mathop{\mathbb{E}}_{#1}[Y_{k}]$ we denote the conditional expectation of $Y_{k}$ wrt “the past”.
+where with $\mathop{\mathbb{E}}_{#1}[Y_{k}]$ we denote the conditional expectation of $Y_{k}$ w.r.t. “the past”.
 Let $X_{k}=Y_{k}-Y_{k-1}$ for $k=1,2,\dots$, which implies $\{X_{k}\}_{k}$ is a martingale difference sequence for $\{Y_{k}\}_{k}$, i.e.,
@@
 while, for $i\neq j$,
-$\mathop{\mathbb{E}}_{#1}[\mathbbm{1}\{X_{t-1}^{(i)}=u\}\mathbbm{1}\{X_{t}^{(j)}=v\}]=\mathbbm{1}\{X_{t-1}^{(i)}=u\}\sum_{w\in V}\mathbbm{1}\{X_{t-1}^{(j)}=w\}P(w,v)$
+$\mathop{\mathbb{E}}_{#1}[\mathbbm{1}\{X_{t-1}^{(i)}=u\}\mathbbm{1}\{X_{t}^{(j)}=v\}]$
+$\quad=\mathbbm{1}\{X_{t-1}^{(i)}=u\}\sum_{w\in V}\mathbbm{1}\{X_{t-1}^{(j)}=w\}P(w,v).$
 Therefore,
@@
 which implies that $Y_{t}$ is a martingale difference sequence.
-Let $\widetilde{N}(u,v)=\sum_{t=0}^{T-1}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\}$. Then, for any $u,v\in V$,
+Let $\widetilde{N}(u,v)=\sum_{t=1}^{T}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\}$. Then, for any $u,v\in V$,
 $\sum_{t=1}^{T}Y_{t}(u,v)$ =|$=\left|\{i,j\in[k],t\in[T]\colon X^{(i)}_{t-1}=u,X^{(j)}_{t}=v\}\right|$
 |$-\left|\{i\in[k],0\leq t<T\colon X^{(i)}_{t}=u\}\right|P(u,v)$
-$-\sum_{w\in V}\sum_{t=0}^{T-1}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=w\}P(w,v)$
+$-\sum_{w\in V}\sum_{t=1}^{T}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=w\}P(w,v)$
 $=$ $N(u,v)-N(u)P(u,v)-\sum_{w\in V}\widetilde{N}(u,w)P(w,v).$
@@
 By Lemma 3, with high probability, $\|\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}]\|<\epsilon/10$. Therefore,
-$\|\widetilde{N}P-k(k-1)T\pi\pi^{T}\|$ $<\|\mathop{\mathbb{E}}[\widetilde{N}]P-k(k-1)T\pi\pi^{T}\|+\|(\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}])P\|$
+$\|\widetilde{N}P-k(k-1)T\pi\pi^{T}\|$ $<\|\mathop{\mathbb{E}}[\widetilde{N}]P-k(k-1)T\pi\pi^{T}\|$
+$\quad+\|(\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}])P\|$
 $\leq\|\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}]\|\|P\|$
@@
 $\|Y_{t}\|_{1}=\max_{v\in V}\sum_{u\in V}|Y_{t}(u,v)|$
-$\leq\max_{v\in V}\left\{|\sum_{i\colon X^{(i)}_{t}=v}(1-P(X^{(i)}_{t-1},v))|+k\cdot\sum_{i=1}^{k}|\sum_{j\colon X^{(j)}_{t}=v}(1-P(X^{(j)}_{t-1},v))|\right\}$
+$\leq\max_{v\in V}\left\{\left|\sum_{i\colon X^{(i)}_{t}=v}(1-P(X^{(i)}_{t-1},v))\right|\right.$
+$\quad+k\cdot\sum_{i=1}^{k}\left.\left|\sum_{j\colon X^{(j)}_{t}=v}(1-P(X^{(j)}_{t-1},v))\right|\right\}$
 $\leq k+(k-1)k=k^{2}.$
 For the $\ell_{\infty}$ norm, first notice that
-$\sum_{v\in V}\left|\left(\mathbbm{1}\{X_{t}^{(i)}=v\}-P(u,v)\right)\right|$ $=1-P(u,X_{t}^{(i)})+\sum_{w\neq X_{t}^{(i)}}P(u,w)$
-$=2(1-P(u,X_{t}^{(i)})).$
+$\sum_{v\in V}\left|\left(\mathbbm{1}\{X_{t}^{(i)}=v\}-P(u,v)\right)\right|$
+$\quad=1-P(u,X_{t}^{(i)})+\sum_{w\neq X_{t}^{(i)}}P(u,w)$
+$\quad=2(1-P(u,X_{t}^{(i)})).$
 Therefore,
 $\|Y_{t}\|_{\infty}$ $=\max_{u\in V}\sum_{v\in V}|Y_{t}(u,v)|$
-$\leq\max_{u\in V}\left\{2\sum_{i=1}^{k}(1-P(u,X_{t}^{(i)}))+\sum_{i=1}^{k}\cdot|\sum_{j\neq i}2(1-P(X^{(j)}_{t-1},v))|\right\}$
+$\leq\max_{u\in V}\left\{2\sum_{i=1}^{k}(1-P(u,X_{t}^{(i)}))\right.$
+$\quad+\left.\sum_{i=1}^{k}\left|\sum_{j\neq i}2(1-P(X^{(j)}_{t-1},v))\right|\right\}$
 $\leq 2k+2(k-1)k=2k^{2}.$
 Hence,
-$\|Y_{t}\|\leq\sqrt{\|Y_{t}\|_{1},\|Y_{t}\|_{\infty}}\leq\sqrt{2}k^{2}.$
+$\|Y_{t}\|\leq\sqrt{\|Y_{t}\|_{1}\cdot\|Y_{t}\|_{\infty}}\leq\sqrt{2}k^{2}.$
 The following lemma provides a bound on $\sigma^{2}$ as required by Theorem 4.
@@
 $=\sum_{w\in V}\sum_{i,i^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}\left(\mathbbm{1}\{X_{j}^{(i)}=u\}-P(w,u)\right)$
-$\qquad\qquad\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=v\}-P(w,v)\right)$
+$\qquad\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=v\}-P(w,v)\right)$
 $+\sum_{w\in V}\sum_{i=1}^{k}\sum_{i^{\prime}\neq\ell^{\prime}}\mathbbm{1}\{X_{j-1}^{(i)}=X_{j-1}^{(i^{\prime})}=w\}\cdot$
@@
 $=\sum_{w\in V}\sum_{i=1}^{k}\mathbb{E}_{j-1}\Bigl{[}\mathbbm{1}\{X_{j-1}^{(i)}=w\}\left(\mathbbm{1}\{X_{j}^{(i)}=u\}-P(w,u)\right)$
-$\qquad\left(\mathbbm{1}\{X_{j}^{(i)}=v\}-P(w,v)\right)\Bigr{]}$
+$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(i)}=v\}-P(w,v)\right)\Bigr{]}$
 $=\sum_{i=1}^{k}\sum_{w\in V}\mathbbm{1}\{X_{j-1}^{(i)}=w\}\Bigl{(}\mathbbm{1}\{u=v\}P(w,u)(1-P(w,u))$
@@
 $\quad-\mathbbm{1}\{u\neq v\}\sum_{w\in V}N(w)P(w,u)P(w,v).$
-Moreover, by Lemma 3, with high probability, we have that $N(u)\leq 2kT\pi(u)$ for any $u\in V$. Therefore,
+By Lemma 3, with high probability, we have that $N(u)\leq 2kT\pi(u)$ for any $u\in V$. Therefore,
 $\left\|\sum_{j=1}^{T}A_{t}\right\|_{1}$ $\leq 2\max_{u\in V}\sum_{w\in V}N(w)P(w,u)$
@@
 Since $\min_{u}D_{N}(u,u)\geq(9/10)\cdot Tk\pi_{\star}$ (again, with high probability), we have that $\|M-(P+(k-1)\Pi)\|\leq\epsilon$.
-8 Acknowledgements
-All authors were partially supported by the Defence Science and Technology Laboratory.
-[…]
+This ends the proof of 1.